# SILMM: Self-Improving Large Multimodal Models for Compositional Text-to-Image Generation

### v1

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated impressive capabilities in multimodal understanding and generation, pushing forward advancements in text-to-image generation. However, achieving accurate text-image alignment for LMMs, particularly in compositional scenarios, remains challenging. Existing approaches, such as layout planning for multi-step generation and learning3 from human feedback or AI feedback, depend heavily on prompt engineering, costly human annotations, and continual upgrading, limiting flexibility and scalability. In this work, we introduce a model-agnostic iterative self-improvement framework (SILMM) that can enable LMMs to provide helpful and scalable self-feedback and optimize text-image alignment via Direct Preference Optimization (DPO). DPO can readily applied to LMMs that use discrete visual tokens as intermediate image representations; while it is less suitable for LMMs with continuous visual features, as obtaining generation probabilities is challenging. To adapt SILMM to LMMs with continuous features, we propose a diversity mechanism to obtain diverse representations and a kernel-based continuous DPO for alignment. Extensive experiments on three compositional text-to-image generation benchmarks validate the effectiveness and superiority of SILMM, showing improvements exceeding 30% on T2I-CompBench++ and around 20% on DPG-Bench. The code and models are available at https://silmm.github.io/.
@@ 1 Introduction
Large Multimodal Models (LMMs) are advancing rapidly, surpassing Large Language Models (LLMs) by embracing multimodal capabilities for multimodal content perception, understanding, and generation [61, 19]. In particular, LMMs demonstrate promising abilities in interpreting user input prompts for text-to-image generation (T2I) [52, 54], producing vivid and photorealistic images. However, as shown in Fig. 1(a), achieving precise text-image alignment between generated images and complex prompts remains challenging, especially for compositional prompts involving multiple objects, attributes, counting, and complex relationships [46, 16, 6].
@@
To enhance text-image alignment, existing work falls into two primary research lines. One line focuses on decomposing the T2I task into multiple stages. For example, some methods perform layout planning before generating the image [17, 35, 72]; while some split the image into sections for multi-step generation via multi-agent collaboration [67, 44]. However, these methods depend on extensive multi-step prompt engineering, which risks error accumulation. The second research line emphasizes learning from human feedback (RLHF [41]) to improve text-image alignment [33, 70, 31, 15, 64], or using automatic AI feedback (RLAIF) from strong evaluation approaches or reward models [3, 74]. Nevertheless, it is labor-intensive and costly to obtain extensive high-quality human feedback, which is also often required to train external reward models [5]. Additionally, as LMMs evolve, the external evaluation approaches and reward models may require continual upgrading [42, 74].
@@ 4.1 Experimental Setup
Base Model Settings. We implement our method on DreamLLM (continuous LMM) [13] and SEED-LLaMA (discrete LMM) [19] for all experiments. Details on DPO training are provided in App. 9.
@@
Benchmarks. We evaluate our method on three text-to-image alignment benchmarks and follow their default settings. T2I-CompBench++ [28] consists of 8,000 compositional text prompts organized into 4 main categories: attribute, layout, non-spatial and complex compositions, further divided into 8 subcategories, including color binding, shape binding, texture binding, 2D/3D-spatial relationships, non-spatial relationships, numeracy, and complex compositions. TIFA [27] uses pre-generated question-answer pairs and a VQA model to evaluate generation results based on 4,081 diverse text prompts and 25,829 questions across 12 categories. DPG-Bench [26] comprises 1,065 densely descriptive prompts with an average token length of 83.91, presenting more complex scenarios with varied objects and rich adjectives.
@@ 4.2 Performance Comparison
Table 1: Performance comparison and improvement of the proposed method for compositional text-to-image generation on T2I-CompBench++ [28], DPG-Bench [26], and TIFA [27]. Alignment scores are calculated using expert understanding models (e.g., VQA or object detection models) recommended by these benchmarks.
@@ 4.4 Qualitative Results
To illustrate the improvements achieved by SILMM, Fig. 8 presents examples generated by SEED-LLaMA, DreamLLM, and SILMM, in Fig. 8 using prompts from T2I-CompBench++. These results showcase the effectiveness of the proposed self-improvement framework across extensive compositional scenarios.
@@ 5 Conclusion
In this work, we present a self-improvement approach named SILMM to enhance text-image alignment within LMMs, introducing an iterative model-agnostic framework comprising five stages to enable high-quality self-feedback and alignment learning. For continuous LMMs, we propose a dropout-based strategy to diversify image representations and a continuous DPO method, KC-DPO, for optimizing LMMs with preference representation pairs. Extensive experiments on three compositional T2I benchmarks validate the effectiveness and superiority of our SILMM framework.

### v2

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated impressive capabilities in multimodal understanding and generation, pushing forward advancements in text-to-image generation. However, achieving accurate text-image alignment for LMMs, particularly in compositional scenarios, remains challenging. Existing approaches, such as layout planning for multi-step generation and learning3 from human feedback or AI feedback, depend heavily on prompt engineering, costly human annotations, and continual upgrading, limiting flexibility and scalability. In this work, we introduce a model-agnostic iterative self-improvement framework (SILMM) that can enable LMMs to provide helpful and scalable self-feedback and optimize text-image alignment via Direct Preference Optimization (DPO). DPO can readily applied to LMMs that use discrete visual tokens as intermediate image representations; while it is less suitable for LMMs with continuous visual features, as obtaining generation probabilities is challenging. To adapt SILMM to LMMs with continuous features, we propose a diversity mechanism to obtain diverse representations and a kernel-based continuous DPO for alignment. Extensive experiments on three compositional text-to-image generation benchmarks validate the effectiveness and superiority of SILMM, showing improvements exceeding 30% on T2I-CompBench++ and around 20% on DPG-Bench. The code is available at https://silmm.github.io/.
@@ 1 Introduction
Large Multimodal Models (LMMs) are advancing rapidly, surpassing Large Language Models (LLMs) by embracing multimodal capabilities for multimodal content perception, understanding [48, 34, 35], and generation [64, 19]. In particular, LMMs demonstrate promising abilities in interpreting user input prompts for text-to-image generation (T2I) [55, 57], producing vivid and photorealistic images. However, as shown in Fig. 1(a), achieving precise text-image alignment between generated images and complex prompts remains challenging, especially for compositional prompts involving multiple objects, attributes, counting, and complex relationships [49, 16, 6].
@@
To enhance text-image alignment, existing work falls into two primary research lines. One line focuses on decomposing the T2I task into multiple stages. For example, some methods perform layout planning before generating the image [17, 37, 75]; while some split the image into sections for multi-step generation via multi-agent collaboration [70, 46]. However, these methods depend on extensive multi-step prompt engineering, which risks error accumulation. The second research line emphasizes learning from human feedback (RLHF [43]) to improve text-image alignment [33, 73, 31, 15, 67], or using AI feedback (RLAIF) from strong evaluation approaches or reward models [3, 77]. Nevertheless, it is labor-intensive and costly to obtain extensive high-quality human feedback, which is also often required to train external reward models [5]. Additionally, as LMMs evolve, the external evaluation approaches and reward models may require continual upgrading [44, 77].
@@ 4.1 Experimental Setup
Base Model Settings. We implement our method on DreamLLM (continuous LMM) [13] and SEED-LLaMA (discrete LMM) [19] for all experiments. We also apply our method to Emu-3 [68], the recent state-of-the-art discrete LMM. Details on DPO training are provided in App. 9.
@@
Benchmarks. We evaluate our method on three text-to-image alignment benchmarks and follow their default settings. T2I-CompBench++ [28] consists of 8,000 compositional text prompts organized into 4 main categories: attribute, layout, non-spatial, and complex compositions, further divided into 8 subcategories, including color, binding, binding, 2D/3D-spatial relationships, non-spatial relationships, numeracy, and complex compositions. TIFA [27] uses pre-generated question-answer pairs and a VQA model to evaluate generation results based on 4,081 diverse text prompts and 25,829 questions across 12 categories. DPG-Bench [26] comprises 1,065 densely descriptive prompts with an average token length of 83.91, presenting more complex scenarios with varied objects and rich adjectives.
@@ 4.2 Performance Comparison
Table 1: Performance comparison and improvement of the proposed method for compositional text-to-image generation on T2I-CompBench++ [28], DPG-Bench [26], and TIFA [27]. Alignment scores are calculated using expert understanding models (e.g., VQA or object detection models) recommended by these benchmarks. Prompt rewriting in Emu3 [68] was not used for fair comparison.
@@ 4.4 Qualitative Results
To illustrate the improvements achieved by SILMM, Fig. 8 presents examples generated by SEED-LLaMA, DreamLLM, and SILMM, in Fig. 8 using prompts from T2I-CompBench++. These results showcase the effectiveness of SILMM across extensive compositional scenarios.
@@ 5 Conclusion
In this work, we present a self-improvement approach named SILMM to enhance text-image alignment within LMMs, introducing an iterative model-agnostic framework comprising five stages to enable high-quality self-feedback and alignment learning. For continuous LMMs, we propose a dropout-based strategy to diversify image representations and a continuous DPO method, KC-DPO, for optimizing LMMs with preference representation pairs. Extensive experiments validate the effectiveness and superiority of our SILMM framework.